OpenStudy (jchick):

Select the system of equations that corresponds to the given graph. graph of two lines. One line goes through the points 0 comma 2 and 1 comma 0. The other line goes through the points 0 comma 2 and negative 1 comma negative 1

OpenStudy (jchick):

a. 4x + y = 6 6x + 3y = 6 b. −3x + y = 2 4x + 2y = 4 c. 2x + y = 4 6x + 3y = 12 d. 4x + 2y = 6 6x + 3y = 12

OpenStudy (jchick):

Answer is B

OpenStudy (jchick):

right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

its c

OpenStudy (jchick):

Oh ok I see where I went wrong thanks!

OpenStudy (anonymous):

:D